OpenStudy (sette):

What is the mass of 1.7 × 1023 atoms of zinc (Zn)?

OpenStudy (sette):

Please Explain! i dont understand

OpenStudy (turingtest):

you need to do a conversion. From the periodic table you can get the molar mass of Zn, which is the number of grams of zinc in one mole of zinc (6.022x10^23 atoms) from that you can find what fraction of a mole of atoms 1.7 x 10^23 is, and use that fraction to determine the weight of the sample.

OpenStudy (sette):

molar mass of Zinc is 65.4

OpenStudy (turingtest):

#atoms x (mass/mole of atoms) a mole is Avagadro's number, 6.022 x 10^23, so #atoms x (mass/6.022 x 10^23 atoms)=mass

OpenStudy (turingtest):

what fraction of a mole is 1.7 x 10^23 ?

OpenStudy (sette):

where do I get the number of atoms from ?

OpenStudy (turingtest):

you are given that you have 1.7 x 10^23 atoms of Zn in the problem

OpenStudy (sette):

1.7x10^23 x (65.4/6.022x10^23) = mass?

OpenStudy (turingtest):

right :) notice how the units cancel

OpenStudy (sette):

18.46 !

OpenStudy (turingtest):

\[1.7\times10^{23}\cancel{\text{atoms}}\times\left(\frac{65.4\text{grams}}{6.022\times10^{23}\cancel{\text{atoms}}}\right)\approx18.46\text{grams}\]it is important to keep track of the units.

OpenStudy (turingtest):

good job :)

OpenStudy (sette):

wow . thank you very much !

OpenStudy (turingtest):

welcome :)

thomaster (thomaster):

This diagram can be really helpful for these kinda problem :)

OpenStudy (sette):

Thanks !